# The Internal Model Principle of Time-Varying Optimization

### v2

@@ 3.1 Fundamental results
We say that a mapping $H_{c}:\Theta\to{}^{n}$ zeros the gradient at the point $\theta\in\Theta$ if
00 0 $=\nabla_{x}f(H_{c}(\theta),\theta).$ (16)
@@
A point $\theta_{\omega}\in\Theta$ is a limit point with respect to the initialization $\theta_{\circ}\in\Theta$ if there exists a sequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ with $t_{i}\to\infty$ as ${i\to\infty}$ such that the exosystem (2) with $\theta(0)=\theta_{\circ}$ satisfies $\theta({t_{i}})\to\theta_{\omega}$ as ${i\to\infty}$. Let $\Omega(\theta_{\circ})$ denote the set of all limit points (i.e., for all sequences $t_{i}$) of (2) with respect to the initialization $\theta_{\circ}\in\Theta.$ Given $\Theta_{\circ}\subseteq\Theta,$ the set $\Omega(\Theta_{\circ}):=\cup_{\theta_{\circ}\in\Theta_{\circ}}\Omega(\theta_{\circ})$ is called the limit set with respect to initializations in $\Theta_{\circ}$ [49$\square$
@@
Answer to Problem 2. When a parameter-feedback algorithm that achieves exact tracking exists, the minimal knowledge needed to design such an algorithm is (O2). Notice that assumption (E1) is implicitly required to execute the algorithm (14).
@@
On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig. 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (16). This graphical observation can be formalized with the aid of the implicit function theorem [47], as described next. Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow{}^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied. $\square$
@@
Answer to Problem 3. When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge (O2) is to first determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$ and then find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set, in which case the parameter-feedback algorithm is given by (14).
@@
with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
@@
For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique. Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin. Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$. Note that, by substituting into (15), we have ${y(t)=RH_{c}\theta(t)+Q\theta(t)=0}$ for all times $t\in\mathbb{R}_{\geq 0}.$ Namely, the gradient is identically zero. This property holds because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin). In fact, this feature holds for the any optimization problem for which the gradient of the cost admits a mapping that zeros the gradient globally. A set of sufficient conditions to check for this property are given in Theorem 3. $\square$
@@ 3.2 Sufficient conditions for the existence of a parameter-feedback algorithm
Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
@@
Then, the result follows from the implicit function theorem (see, e.g., [50, Thm. 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$. ∎
@@
Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin. The following result provides a sufficient condition for global convergence.
@@
    the loss function $f$ is $C^{1}$ on ×${}^{n}\times\Theta$,
@@
    the Hessian $\nabla^{2}_{xx}f(x,\theta)$ is positive definite on ×${}^{n}\times\Theta$,
@@
    the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on ×${}^{n}\times\Theta.$
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$square
@@
Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on ×${}^{n}\times\Theta$, the mapping $x\mapsto F(x,\theta)$ is $C^{2}$ on n for all $\theta\in\Theta$,
$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]\neq 0,\quad\forall(x,\theta)\in{}^{n}\times\Theta,$
and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$. Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [50, Thm. 2] (see also [51, Thm. 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows. ∎
Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$
@@ 4 The dynamic gradient-feedback problem
The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [52, Cor. 3.4]. Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (and, as a consequence, the critical points of (1)), and thus they can be removed from the optimization problem (1) without altering its critical points. In other words, Assumption 3 ensures that (2) does not contain redundancies.
@@ 4.1 Existence and characterization of dynamic gradient feedback algorithms
00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
@@
(If). We now prove that (23) implies ${\lim_{t\rightarrow\infty}y(t)=0}$. It follows from (23) and Theorem 10 that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12). Moreover, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1. ∎
@@
is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [52, Sec. 3].
@@
We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact asymptotic tracking. Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm (see (7)). $\square$
@@
Recall that the gradient-flow algorithm (4) can be viewed as an instance of (10a) with $F_{c}(z,y)$ and $G_{c}(z)$ as in Remark 7. By direct substitution into (23), it is immediate to see that this algorithm satisfies (23a) with $s(\theta)=0$ and $\sigma(\theta)$ arbitrary. Since $s(\theta)=0$ describes the internal model of a constant signal, it follows from Theorem 4 that these algorithms achieve exact asymptotic tracking if and only if $\theta(t)$ is a constant signal. This observation is in line with the inexact convergence properties of these algorithms typically given for these algorithms [48, 26, 14$\square$
Since by, Theorem 4, the problem of exact asymptotic tracking can be reduced to that of finding an optimization algorithm that is algebraically equivalent to the exosystem and a mapping that zeros the gradient, we have the following.
Answer to Problem 2.2. When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2). Notice that assumption (O1) is implicitly required to execute the algorithm (10).
By comparison with Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
@@
An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$square
@@
Suppose Assumptions 13 hold. Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations. $\square$
@@
(If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$. First, notice that by Assumption 3, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$. Consider the algorithm (10) with $n_{c}=p$ and [^6] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
@@
where $H_{c}(z)$ is as in (16). The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22). The claim thus follows by Theorem 4 with $\sigma$ the identity operator on $\Theta$. ∎
Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for the existence of a parameter-feedback algorithm (cf. Section 3) are identical. This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (27), the dynamic state of the controller $z(t)$ acts as an alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (26)).

### v3

@@ 3.1 Fundamental results
We say that a mapping $H_{c}:\Theta\to\real^{n}$ zeros the gradient at the point $\theta\in\Theta$ if
$0$ $=\nabla_{x}f(H_{c}(\theta),\theta).$ (16)
@@
[…]
@@
Answer to Problem 2: When a parameter-feedback algorithm exists, the minimal knowledge needed to design such an algorithm is (O2). Notice that assumption (E1) is implicitly required for (14) to be implementable.
@@
On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig. 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (16). This graphical observation can be formalized with the aid of the implicit function theorem [48], as described next. Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow\real^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied. $\square$
@@
Answer to Problem 3: When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge is given by: 1) determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$; 2) find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set; and, 3) the parameter-feedback algorithm is given by (14).
@@
with matrices $R\in\mathbb{S}^{n}$ and $Q\in\real^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
@@
For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique. Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in\real^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin. Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$. Note that, by substituting into (15), we have $y(t)=RH_{c}\theta(t)+Q\theta(t)=0$} for all times $t\in\mathbb{R}_{\geq 0}.$ Namely, the gradient is identically zero. This property holds because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin). More generally, this feature will hold for any problem for which the gradient of the cost admits a mapping that zeros the gradient globally. A set of sufficient conditions to check for this property are given in Theorem 3. $\square$
@@ 3.2 Sufficient conditions for the existence of a parameter-feedback algorithm
Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
@@
Then, the result follows from the implicit function theorem (see, e.g., [51, Thm. 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$. ∎
@@
Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature; namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin. The following result provides a sufficient condition for global convergence.
@@
    the loss function $f$ is $C^{1}$ on $\real{}^{n}\times\Theta$,
@@
    the Hessian $\nabla^{2}_{xx}f(x,\theta)$ is positive definite on $\real{}^{n}\times\Theta$,
@@
    the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on $\real{}^{n}\times\Theta.$
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient everywhere in $\Theta.$ $\square$
@@
Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on $\real{}^{n}\times\Theta$, the mapping $x\mapsto F(x,\theta)$ is $C^{2}$ on n for all $\theta\in\Theta$,
$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]\neq 0,\quad\forall(x,\theta)\in\real^{n}\times\Theta,$
and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in $\real{}^{n}\times\Theta$. Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [51, Thm. 2] (see also [52, Thm. 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows. ∎
Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient everywhere in $\Theta.$
@@ 4 The dynamic gradient-feedback problem
The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [53, Cor. 3.4]. Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (at $(x_{\circ}^{\star},0)$) and, as a consequence, the critical points of (1); therefore, they can be removed from the optimization problem (1) without altering its critical points. In other words, Assumption 3 ensures that (2) does not contain redundancies.
@@ 4.1 Existence and characterization of dynamic gradient feedback algorithms
$0$ $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
@@
(If). We now prove that (23) implies $\lim_{t\rightarrow\infty}y(t)=0$}. It follows from (23) that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12). Hence, by Theorem 10, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1. ∎
@@
is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [53, Sec. 3].
@@
We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact tracking. Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm (see (7)). $\square$
@@
[…]
Since Theorem 4 shows that the problem of exact asymptotic tracking can be reduced to finding an optimization algorithm that is algebraically equivalent to the exosystem together with a mapping that zeros the gradient, we conclude the following.
Answer to Problem 2.2: When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2). Notice that assumption (O1) is implicitly required to execute the algorithm (10).
By comparison with the Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
@@
An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$ $\square$
@@
Suppose Assumptions 13 hold. Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to\real^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations. $\square$
@@
(If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$. First, notice that by Assumption 3, there exists a matrix $L$ such that SLQS-LQ has eigenvalues in $\mathbb{C}^{-}$. Consider the algorithm (10) with $n_{c}=p$ and [^6] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
@@
where $H_{c}(z)$ is as in (16). The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by SLQS-LQ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since SLQS-LQ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22). The claim thus follows by Theorem 4 with $\sigma$ the identity operator on $\Theta$. ∎
Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for a parameter-feedback algorithm (cf. Section 3) are identical. This should not be surprising: on the one hand, a parameter-feedback algorithm assumes access to $\theta(t)$; on the other hand, a gradient-feedback algorithm must infer $\theta(t)$ through $y(t)$, but (see (27)) the dynamic state of the controller $z(t)$ acts as a copy (possibly in different coordinates) of $\theta(t)$. In this sense, both algorithm classes rely on essentially the same information, and it is therefore natural that the conditions for their existence coincide.